# EcoServe: Designing Carbon-Aware AI Inference Systems

### v1

@@ VI-A End-to-end Evaluation
  • Reduce can yield around 16-42% carbon savings. For leaner GPU offerings like T4, the savings are less than higher-end GPUs since the host is designed to scale with the GPU memory capacity.
  • Right sizing (RS) enables around 15-39% better carbon, leveraging the SLO slacks for various workloads on TPOT and TTFT. The benefit is particularly significant in lower carbon intensity regions with higher load variability and higher online workload component, as EcoServe can dynamically adjust its resource allocation to match demand to different workload slices.
  • Optimized CPU reuse strategies can save around 25-30% of carbon for clusters with different percentages of offline workload.
@@
Figure 14: Beginning with llama.cpp as a baseline, we optimize CPU attention for Gemma-2b and Gemma-27b GGUF models. EcoServe’s CPU decode optimization improve performance by up to 4.03$\times$ and on average by 1.34$\times$ (across batch-sizes and parallelism dimensions) on an Intel Sapphire Rapids 56 core CPU.
VI-A1 Performance Analysis
From Figure 13, we demonstrate that the performance of EcoServe variants is close to Perf-opt within 3 percent, except for the reuse case. The reason for a worse TPOT (around 5%) is because of the offline workloads, where the reuse strategy prioritize the decoding of tokens with longer context length to be offloaded to CPU. This is not perf-optimal when there are further GPU capacities to use for serving offline workloads.
EcoServe (RS) provides better improvements with TPOT metric than TTFT metric in the presence of SLO slack. This is because of the longer prompt and higher load, when we are resizing the allocation, the unavailability will cause the prompting scheduler to temporarily queue the unserved requests’ prompts, and the queuing delay increases TTFT. Furthermore, by design, our Reuse mostly addresses the decoding servers’ capacity planning with CPUs, hence the performance on TTFT on the prompting servers is untapped. EcoServe (Reduce) incurs a little overhead on TPOT (¡5%); but energy-opt allocator is worse in performance through aggressively choosing energy-opt hardware, ignoring the embodied carbon.
VI-A2 Breakdown Analysis
@@
For online workloads, Reduce (online resource) allows us to save approximately 17-40% of the carbon footprint by reducing DRAM and SSD capacity. When moving to Reuse, online workloads do not use CPU resources, maintaining consistent savings. The RS (online) strategy further increases savings by employing workload-aware load balancers and resource allocators, limiting the overprovisioning of high-end hardware.
@@
VI-B Reuse Strategy Evaluation
@@
We evaluate the two design aspects of the reuse strategy: first, optimized CPU threading and tiling implementation for GEMM and long-context decoding. Second, workload- and load-aware CPU reuse strategy. Figure  14 shows the speedup with various workload sizes with EcoServe’s optimized reuse strategy, compared with state-of-the-art CPU inference engine baseline (llama.cpp) on Gemma 2/27B. The evaluation varies the core counts (blue as 56 and black as 112) to show the generality of the approach. Through choosing the right tiling and parallelism strategy guided by performance modeling, we improve decoding latency, which directly translate to higher throughput and carbon savings of 1.34x on average.
@@
Figure 15: Decoding throughput (Left), operational (Middle) and embodied carbon (Right), for various models and workload sizes, comparison with naive llama.cpp vs. optimized CPU reuse strategy, normalized to A100 GPU performance under maximum throughput. Carbon reduction is 12, 26% for Gemma-27B, and 40%, 66% for Llama-8B respectively under short and longer workload.
@@
Figure 15 analyzes the detailed embodied and operational carbon breakdowns compared with serving on GPUs for EcoServe’s reuse, showing the benfit of workload- and load-aware optimization. Given the CPU’s lack of energy proportionality, the added operational power is relatively smaller, but still results in operational carbon cost with longer execution time. The free lunch from a 56-core SPR machine CPU attached to the A100 enables us to have 0.53-2.29$\times$ of the throughput of the GPU (Blue bars). Although CPU generally has worse energy efficiency, especially for larger models when both CPU and GPU are compute bound (Gemma, Orange bars) EcoServe’s CPU decode can achieve up to 66% relative embodied carbon savings compared with a naive llama.cpp (CPU-Naive) strategy as 48%. On average, EcoServe can achieve 3.51x embodied carbon savings compared with non optimized baseline across the above workloads. More importantly, without EcoServe’s reuse strategy (CPU-Opt), the embodied carbon for llama.cpp (CPU-naive) is worse than non-reuse (GPU) for Gemma-27B model and Llama-8B model short context cases. This is due to the fact that under longer context and smaller model dimension, the gap between CPU and GPU throughput is larger due to batching effect, and the carbon saving is larger in an iso-throughput setting.
The total carbon benefit comes from the net effect of embodied and operational carbon with Reuse. Although the operational carbon is not better than GPU, by leveraging the unsused CPU cores, EcoServe improves both in cost and embodied carbon; fewer GPUs need to be over-provisioned. Furthermore, from throughput perspective, EcoServe demonstrates more benefits for long context decode with customized CPU kernel optimization, and we can enable flexible carbon reduction with runtime scheduling based on workload sizes and demands.
@@
Figure 16: Relative carbon and energy improvements of rightsizing GPUs on Gemma-27B on homogeneous GPUs (i.e., H100, A100, L4), and cost-optimized provisioning (Melange). EcoServe’s Right-sizing strategy exploits LLM phases to reduce overall energy and carbon under TPOT = 100 ms (online) and TTFT = 24hr offline constraints.
VI-C Right-sizing Strategy Evaluation
For right-sizing, we evaluate our system on Gemma-27B with various load and workload against the following two types of baselines:
  • Melange: Cost-optimal GPU resource allocation.
@@ VI-C Right-sizing Strategy Evaluation
  • Single hardware: pick the single hardware and its replica numbers to satisfy workload demands that disregard SLO slacks and the energy/carbon implication for hardware choices.
@@
Figure 16 shows the final outcomes for right-sizing strategy given different workload types and load of the system. Each request generator runs for 5 minutes length traces, and we discard the first and last 5% of the statistics for warming up effects. The length of the workloads are randomly sampled from the open-sourced traces with both long and short prompts.
@@
For online setting (left), under lower request rate, EcoServe consistently outperforms Melange by over 2.56x for carbon and 2.25x for energy. The improvement owes to separate the resource allocation decisions for LLM phases; finer-grained allocation enables better utilization and energy. As an example, for request rate equal to 1, EcoServe will choose the most carbon efficient hardware based on the workload size; e.g. for medium-sized workload it chooses L4 for decoding and A100 for prompting; for long prompt it chooses A100 for decoding and H100 for prompting. However, Melange sticks to the best perf-per-cost hardware for both phases for each workload slice, and the quantization effect for available hardware capacity introduces more over-provisioned resources for decode phase, which negatively impacts embodied and operational carbon.
@@
Furthermore, for offline setting (offline), EcoServe consistently wins in both energy and carbon up to 1.6$\times$ with the next best baseline, as it considers both dimensions in the optimization objective. Under the carbon intensity of US-Central / South, operational carbon is around 0.64x of the embodied, making EcoServe a desirable solution for balancing both energy and embodied carbon.
VI-D Recycle Strategy Evaluation
Figure  17 shows the absolute carbon cost (top) and relative carbon savings (bottom) across years with a fixed (baseline) and asymmetric (EcoServe) upgrade strategy. The baseline server configuration assumes an embodied carbon of 800 kgCO2e for the host and 120 kgCO2e for the GPU. Additionally, the operational carbon emissions per year for the server are set at 600 kgCO2e. We assume that hardware GPU energy efficiency doubles every 3.5 years [63]. In the baseline fixed upgrade strategy, both the host systems and GPU accelerators have the same 4-year upgrade cycle. EcoServe, however, proposes an asymmetric upgrade strategy where host systems are upgraded less frequently (every 10 years) while GPUs are upgraded more often (every 3 years) to better align with their energy efficiency improvements.
@@ VI-D Recycle Strategy Evaluation
Figure 17: Trends in annual embodied and operational carbon based on EcoServe’s assymetric recycling of CPUs (10 years) and GPUs (3 years). Baseline is a fixed schedule for both CPU and GPU (4 years). Assuming energy efficiency doubles every 3.5 year [63].
VII Related Work
@@
Existing CPU Inference LLM inference on CPUs, as demonstrated by frameworks like Llama.cpp and Gemma.cpp, has garnered significant attention for enabling cost-effective deployment [25, 26]. These systems employ optimizations such as quantization and memory-mapped weights to run models with limited resources. Recent studies also focus on improving CPU-aware inference efficiency through sparsity-aware techniques, quantization optimizations, offloading, and library enhancements (e.g., OpenBLAS [18], oneDNN [51], XNNPACK [56]), some of which are implemented in llama.cpp [25]. We built our work on top of llama.cpp and come up with new solutions for codesigning resource provisioning and workload scheduling.
@@
Heterogeneous LLM on Commodity Hardware: Recently, people have gained more interest in deploying or training foundational models under decentralized or heterogeneous commodity setups [72, 58]. To account for the heterogeneity in the interconnect bandwidth of commodity setup, Mobius [23] introduces a novel pipeline parallelism scheme enabling heterogeneous memory for large-scale model training, while bringing fewer communications than existing systems. While those works mostly focus on model partition and placement, we focus more on the request scheduling and resource provision. Our objective which includes carbon is also fundamentally different from performance-oriented systems.
@@
CPU Memory Offloading Recently, memory offloading has gained popularity as a research area due to the limitations and high costs of GPU HBM. DeepSpeed ZeRO-Inference [7] can run LLMs with offloading. Specifically for MoE models, Huang et al [33] propose to swap experts to CPU memory to reduce the GPU memory consumption. PC-MoE [37] proposes to offload experts under dynamic resource constraints with offline profiling and online scheduling. SiDA [20] dynamically load activated experts and offload inactivated experts according to hash tables. These proposals are mostly focused on MoE models. While helpful to reduce carbon footprint, we consider them to be orthogonal. Furthermore, naively implementing offloading without capacity planning would only cause more embodied and operational carbon due to high PCIE overhead.
@@
Energy-aware Serving and Model Serving: A plethora of work in the space of model serving has emerged, focusing on providing SLOs with scheduling, placement, batching and autoscaling solutions. For example, AlpaServe [40], Clockwork [29], Sheperd [73] and Clipper [14] all are focusing on non-autoregreesive models. However, such general systems often overlook the distinct phases in LLM serving or autoregressive nature of LLMs. InFaaS [57] proposes a way to do model selection and autoscaling. More recently, userve [54] proposed power-aware LLM serving. More recently, Splitwise propose pd disaggregation for optimizing for cost, power, and throughput under SLOs  [53]. These general systems are not carbon-aware. Some newer works like Clover shows the tradeoffs between accuracy and carbon [39]. We show in our experiments that carbon-opt is different from energy-opt, and propose co-design solutions for both resource provision and runtime scheduling, leaning on the insights of asymmetric carbon cost of CPU and GPUs, without hurting model quality.
VIII Conclusion
Our work presents a framework for quantifying AI carbon footprints by integrating operational and embodied emissions, enabling AI sustainability optimization across various system stacks. It lays the groundwork for carbon-aware models, system-hardware co-design, and better hardware refresh strategies. We urge the research community to advance sustainable AI through holistic, quantitative approaches to environmental impact assessment and mitigation.

### v2

@@ 6.1. End-to-end Evaluation
  • Reduce can yield around 12.4-28.6% carbon savings. For leaner GPU offerings like T4, the savings are less than higher-end GPUs since the host is designed to scale with the GPU memory capacity.
  • Right sizing (RS) enables around 15.2-30.3% better carbon, leveraging the SLO slacks for various workloads on TPOT and TTFT. The benefit is particularly significant in lower carbon intensity regions with higher load variability and higher online workload component, as EcoServe can dynamically adjust its resource allocation to match demand to different workload slices.
  • Optimized CPU reuse strategies can save up to 25.4% of carbon for clusters with different percentages of offline workload.
  • Recycle strategy is orthogonal to all the other Rs. It can be independently deployed, yielding around 16.8% savings compared with a homogeneous update baseline (Section 6.5).
@@
The results motivate co-designing the 4Rs for different workloads.
6.1.1. Performance Analysis
Figure 16. Carbon savings and reuse configurations sampled by EcoServe under various workload lengths, SLOs and carbon intensities for a mid-sized model (Llama-70B).
From Figure 15, we demonstrate that the performance of EcoServe variants is close to Perf-opt within 3 percent, except for the reuse case. The reason for a worse TPOT (around 5%) is because of the offline workloads, where the reuse strategy prioritize the decoding of tokens with longer context length to be offloaded to CPU. The KV cache transfer This performance drop can be recovered with better hardware interconnects and other system techniques like batching transfer for smaller models to improve MBU (memory bandwidth utilization).
EcoServe (RS) provides better improvements on TPOT metric than TTFT metric in the presence of SLO slack. This is because of the longer prompt and higher load, when we right-size the allocation periodically, the unavailability will cause the prompt scheduler to queue some requests’ prompts temporarily, which results in the queuing delay. Furthermore, by design, Reuse mostly addresses the decoding servers’ capacity saving, hence the performance on TTFT on the prompting servers is untapped. EcoServe (Reduce) incurs a little overhead on TPOT (¡5%), but the energy-opt allocator is worse in performance through aggressively choosing energy-opt hardware (like L4), ignoring the embodied carbon per performance. Furthermore, Splitwise suffers in carbon and performance because of the limited hardware options, and the gap between provision and scheduling decisions (Section 4.1.2). Further analysis on a specific workload is shown in Figure 17.
6.1.2. Breakdown Analysis.
@@ 6.1.1. Performance Analysis
For online workloads, Reduce (online resource) allows us to save approximately 17-40% of the carbon footprint by reducing DRAM and SSD capacity. When moving to Reuse, online workloads seldom use CPU resources, maintaining consistent savings. The RS (online) strategy further increases savings by employing workload-aware load balancers and resource allocators, limiting the overprovisioning of high-end hardware.
@@
6.2. Sensitivity and Scalability Study
6.2.1. Varying Carbon Intensity.
To show the robustness of EcoServe’s carbon savings under various geographical regions, models,, and workload patterns with different carbon intensity levels, we conducted a controlled experiment based on Splitwise code traces (for 2 minutes) on the simulation infrastructure for a medium/large model (Llama-70b, Bloom-176b) with different request arrival rates.
We selected three power grids: North Central Sweden (Low), California (Mid), and Midcontinent (High). These regions represent low, medium, and high carbon intensity levels, with carbon intensities of 17, 261, and 501 gCO2/kWh, respectively. The total carbon is shown in Figure 17. Notably, EcoServe selects the software provision strategy that is different from SplitWise because the runtime could allocate shorter prompts to the A100. This leads to carbon savings on that trace for 26.5% on average.
Under a 40 H100 power-equivalent setup with long input sequences, Splitwise achieved Pareto-optimal energy-throughput by allocating 35 Prompt and 8 Token machines. However, its JSQ scheduling missed workload-aware co-design opportunities. EcoServe matched throughput with 30 H100s and 17 A100s by adaptively placing longer requests to A100s, reducing both embodied and operational carbon by up to 38% and 36% for large and medium models under lower load.
From Figure 16, we observe that for longer requests, lower carbon intensity level, reuse is most often used. The carbon savings for offline workload is most significant. However, when carbon intensity is higher, we save more with right-sizing and reduce strategy for online workloads.
6.2.2. Scalability of the control plane
cluster size online (low) offline (low) online (high) offline (high)
10 0.281 0.343 0.428 0.523
20 0.377 0.521 0.574 0.789
40 0.398 0.672 0.572 1.030
80 0.508 0.883 0.747 1.272
160 0.585 0.919 0.901 1.315
Table 3. Overhead of control plane over different cluster sizes, workloads and system loads (Unit: second).
To show the scalability of EcoServe’s control plane (ILP solver) with various cluster sizes and request arrival rates (low/high), we measured the time taken for the ILP solver and scheduler to make the right-sizing decision. Table 3 shows the control plane overhead across different cluster sizes and load conditions. The overhead grows sub-linearly with cluster size, increasing by only 2.1-2.7$\times$ when scaling from 10 to 160 nodes. This favorable scaling comes from the ILP formulation’s ability to exploit our carbon and latency performance model - similar workloads on similar hardwares are clustered together. Under high load, offline planning exhibits the highest overhead (1.315s at 160 nodes) due to exploring more complex hardware combinations. However, even in this worst case, the sub-2-second latency remains practical for production deployments where planning decisions typically operate on minute- or hour-level intervals. Online decisions show better scaling (0.585s at 160 nodes) by pruning out irrelevant spaces, focusing on immediate resource availability rather than long-term optimization.
Figure 17. Comparison of Bloom-176B (Left) and Llama-70B (Right) with Splitwise on various carbon intensities (Low, Mid, High) and load on iso-power deployment. EcoServe consistently outperforms Splitwise, and the gap is larger under a lower request rate and higher CI.
6.3. Reuse Strategy Evaluation
@@
We evaluate the two design aspects of the reuse strategy: first, optimized CPU threading and tiling implementation for GEMM and long-context decoding. Second, workload- and load-aware CPU reuse strategy. Figure  18 shows the speedup with various workload sizes with EcoServe’s optimized reuse strategy, compared with state-of-the-art CPU inference engine (llama.cpp) on Gemma 2/27B. The evaluation varies the core counts (56 vs 112) to show the generality of the approach. By choosing the right tiling and parallelism strategy guided by performance modeling, we improve decoding latency, which directly translates to higher throughput and carbon savings of 1.34$\times$ on average.
Figure 18. Beginning with llama.cpp as a baseline, we optimize CPU decoding. EcoServe’s CPU optimization improves performance by up to 4.03$\times$ and on average by 1.34$\times$ (across batch sizes and parallelism dimensions) on an Intel SPR.
@@
Figure 19. Decoding throughput (Left), operational (Middle), and embodied carbon (Right), for various models and workload sizes, comparison with a naive llama.cpp vs. optimized CPU reuse strategy, normalized to A100 GPU performance under maximum throughput. Carbon reduction is 12, 26% for Gemma-27B, and 40%, 66% for Llama-8B, respectively, under short and longer workloads.
@@
Figure 19 analyzes the detailed embodied and operational carbon breakdowns compared with serving on GPUs for EcoServe’s reuse, showing the benefit of workload- and load-aware optimization. Given the CPU’s lack of energy proportionality, the added operational power is relatively minor but still results in operational carbon cost with longer execution time. The free lunch from a 56-core SPR machine CPU attached to A100 allows us to achieve 0.53-2.29$\times$ the throughput of the GPU (Blue bars). Although CPU generally has worse energy efficiency, especially for larger models when both CPU and GPU are compute bound (Orange bars), EcoServe’s CPU decode can achieve up to 66% relative embodied carbon savings compared with a naive llama.cpp (CPU-Naive) strategy as 48%. On average, EcoServe can achieve 3.51$\times$ embodied carbon savings compared with a non-optimized baseline across the above workloads. More importantly, without EcoServe’s reuse strategy (CPU-Opt), the embodied carbon for llama.cpp (CPU-naive) is worse than non-reuse (GPU) for Gemma-27B model and Llama-8B model short context cases. This is due to the fact that under a setting with a longer context and smaller model dimensions, the gap between CPU and GPU throughput is larger due to the batching effect, and the carbon saving is more significant in an iso-throughput setting.
The total carbon benefit arises from the net effect of embodied and operational carbon through Reuse. EcoServe enhances cost and embodied carbon by utilizing unused CPU cores, reducing the need for over-provisioned GPUs. Additionally, it offers advantages in long-context decoding through customized CPU kernel optimization and allows flexible carbon reduction via runtime scheduling based on workload sizes. However, it may not work well for shorter-context larger-model decoding in terms of operational carbon. The strategy can be tailored according to the region’s carbon intensity. In areas with high carbon intensity (CI), operational carbon is a primary concern, so we should minimize CPU offloading. Conversely, in regions with low CI, embodied carbon plays a larger role, allowing us to increase capacity through CPU offloading (Figure LABEL:fig:ablation).
@@
Figure 20. Relative carbon and energy improvements of rightsizing GPUs on Gemma-27B on homogeneous GPUs (i.e., H100, A100, L4), and cost-optimized provisioning (Melange). EcoServe’s Right-sizing strategy exploits LLM phases to reduce overall energy and carbon under TPOT = 100 ms (online) and TTFT = 24hr offline constraints.
6.4. Right-sizing Strategy Evaluation
@@
For right-sizing, we evaluate our system on Gemma-27B with various load and workload against the following two types of baselines: 1) Melange: Cost-optimal GPU resource allocation. 2) Single hardware: pick the single hardware and its replica numbers to satisfy workload demands that disregard SLO slacks and the energy/carbon implication for hardware choices.
@@
Figure 20 shows the final outcomes for right-sizing strategy given different workload types and loads of the system. Each request generator runs for 5 minutes length traces, and we discard the first and last 5% of the statistics for warming up effects. The length of the workloads is randomly sampled from the open-sourced traces with both long and short prompts.
@@
For the online setting (left), under a lower request rate, EcoServe consistently outperforms Melange by over 2.56$\times$ for carbon and 2.25$\times$ for energy. The improvement comes from separating resource allocation decisions for LLM phases; finer-grained allocation enables better utilization and energy. As an example, for request rate of 1, EcoServe will choose the most carbon-efficient hardware based on the workload size; e.g. for medium-sized workload it chooses L4 and A100 for decoding and prompting, respectively; and A100 and H100 for long prompt. However, Melange sticks to the best perf-per-cost hardware for both phases for each workload slice, and the quantization effect for available hardware capacity introduces more over-provisioned resources for decode phase, which negatively impacts embodied and operational carbon.
@@
Furthermore, for offline setting (offline), EcoServe consistently wins in both energy and carbon up to 1.6$\times$ with the next best baseline, as it considers both dimensions in the optimization objective. Under the carbon intensity of US-Central / South, operational carbon is around 0.64$\times$ of the embodied, making EcoServe a desirable solution for balancing both energy and embodied carbon.
6.5. Recycle Strategy Evaluation
Figure 21 shows absolute carbon cost (top) and relative savings (bottom) for fixed (baseline) vs. asymmetric (EcoServe) upgrades. The baseline assumes 800 kgCO2e for the host, 120 kgCO2e for the GPU, and 600 kgCO2e in yearly operational emissions. GPU energy efficiency is estimated to double every 3.5 years on average (sun2019summarizing, ). The baseline upgrades both host and GPU every 4 years, while EcoServe upgrades hosts every 9 years and GPUs every 3 years to better match efficiency gains.
@@ 6.5. Recycle Strategy Evaluation
Figure 21. Trends in annual embodied and operational carbon based on EcoServe’s asymmetric recycling of CPUs (9 years) and GPUs (3 years). Baseline is a fixed schedule for both CPU and GPU (4 years).
7. Related Work
@@
Accelerating CPU Inference: LLM inference on CPUs, as demonstrated by frameworks like Llama.cpp and Gemma.cpp, has garnered significant attention for enabling cost-effective deployment (llama_cpp, ; gemma_cpp, ). These systems employ optimizations such as quantization and memory-mapped weights to run models with limited resources. Recent studies also focus on improving CPU-aware inference efficiency through sparsity-aware techniques, quantization optimizations, offloading, and library enhancements (e.g., OpenBLAS (openblas, ), oneDNN (onednn, ), XNNPACK (xnnpack, )), some of which are implemented in llama.cpp (llama_cpp, ). We built our work on top of llama.cpp and come up with new solutions for codesigning resource provisioning and workload scheduling.
@@
Heterogeneous LLM on Commodity Hardware: Recently, people have gained more interest in deploying or training foundational models under decentralized or heterogeneous commodity setups (yuan2022decentralized, ; scao2022bloom, ). To account for the heterogeneity in the interconnect bandwidth of commodity setup, Mobius (Feng2023, ) introduces a novel pipeline parallelism scheme enabling heterogeneous memory for large-scale model training, while bringing fewer communications than existing systems. While those works mostly focus on model partition and placement, we focus more on the request scheduling and resource provision. Our objective which includes carbon is also fundamentally different from performance-oriented systems.
@@
LLM Serving Offloading: Recently, memory or compute offloading has gained popularity as a research area due to the capacity limits and high costs of GPU HBM. Many works offload model weights, activations, KV cache, or computations to the CPU for offline inference that trade latency for throughput. FlexGen (flexgen, ), HeteGen (hetegen, ), PowerInfer (song2023powerinfer, ), and TwinPilots (twinpilots, ) are focused on settings with batch inference of workloads with homogeneous input / output length. InstInfer (instinfer, ) and DeepSpeed ZeRO-Inference (aminabadi2022deepspeed, ) further offloads to SSD, but they did not study the load balance between CPUs and GPUs. FastDecode (fastdecode, ) and NEO (jiang2024neo, ) target online serving. NEO proposes asymmetric GPU-CPU pipelining and load-aware scheduling to balance GPU and CPU loads. Specifically for MoE models, Huang et al (huang_towards_2023, ) propose to swap experts to CPU memory to reduce the GPU memory consumption. PC-MoE (kong_serving_2023, ) proposes to offload experts under dynamic resource constraints with offline profiling and online scheduling. SiDA (Du2023SiDASD, ) dynamically loads activated experts and offloads inactivated experts according to hash tables. None of the proposals considers the resource provisioning aspect and carbon impact of offloading. Furthermore, naively implementing offloading without considering phases, request lengths, and capacity planning would only cause more embodied and operational carbon due to GPU overprovisioning and large PCIe-induced overhead.
@@
Energy-aware Serving and Model Serving: A plethora of work in the space of model serving has emerged, focusing on providing SLOs with scheduling, placement, batching and autoscaling solutions. For example, AlpaServe (li2023alpaserve, ), Clockwork (gujarati2020serving, ), Sheperd (zhang2023shepherd, ) and Clipper (crankshaw2017clipper, ) all are focusing on non-autoregreesive models. However, such general systems often overlook the distinct phases in LLM serving or autoregressive nature of LLMs. InFaaS (romero2021infaas, ) proposes a way to do model selection and autoscaling. More recently, $\mu$serve (qiu2024muserve, ) proposed power-aware LLM serving. Splitwise proposed pd disaggregation to optimize for cost, power, and throughput under SLOs  (patel2023splitwise, ). These general systems are not carbon-aware, nor do they consider the role of CPU. We show in our experiments that carbon-opt is different from energy-opt, and propose co-design solutions for both resource provisioning and runtime scheduling, leaning on the insights of asymmetric carbon cost of CPU and GPUs.
8. Conclusion
Our work introduces a quantitative framework that holistically optimizes AI infrastructure’s carbon footprint. EcoServe demonstrates how hardware-software co-design across software provision, scheduling, and hardware provision can reduce total carbon emissions while maintaining performance. Beyond the specific optimizations presented, this work provides a foundation for carbon-aware system design that we hope will inspire future research in sustainable AI infrastructure.